# Enhancing Diffusion-Based Sampling with Molecular Collective Variables

### v1

@@ E.1 Alanine dipeptide
In this section, we extend the alanine dipeptide sampling comparison from the main text to standard distributional metrics following Liu et al. (2025). Specifically, we report KL divergence for 1-D marginals of torsion angles, as well as $\mathcal{W}_{2}$ and Jensen–Shannon divergence (JSD) for the joint distribution of backbone torsions $(\phi,\psi)$. Results for previous diffusion samplers using internal coordinate representations (bond distances, angles, and torsions) and for models trained in this work using Cartesian coordinates are summarized in Table˜2. We used $10^{6}$ samples to compute $D_{\text{KL}}$ and JSD, and randomly selected $10^{4}$ samples for $\mathcal{W}_{2}$ (with weighted random sampling for WT-ASBS).
@@
Table 2: Distributional metrics for alanine dipeptide. Comparison of diffusion samplers applied in internal or Cartesian coordinate representations of alanine dipeptide sampling. The KL divergence ($D_{\text{KL}}$) of five 1-D torsion angle marginals and the Wasserstein-2 distance ($\mathcal{W}_{2}$) and JSD on joint backbone torsion distribution are reported. Results from Liu et al. (2025). Result from Choi et al. (2025). Reweighted samples are used.
@@
Figure 7: Ramachandran plots for alanine dipeptide. Comparison of backbone torsion distributions $(\phi,\psi)$ in log densities, for reference, ASBS (internal coordinates), ASBS and WT-ASBS (Cartesian coordinates), as in Table˜2. All three methods reproduce the more populated regions reasonably well, but their behavior in less populated regions differs substantially. Negligibly low-density regions are omitted for visual clarity. Taken from Liu et al. (2025).
Table˜2 shows that WT-ASBS performs well on all metrics, with similar $\mathcal{W}_{2}$ performance to baseline ASBS (Cartesian) and slightly worse than reference ASBS (internal coordinates). However, the Ramachandran plots in Figure˜7 reveal that WT-ASBS better reproduces the overall density, while the other two methods either collapse into highly populated regions or leak into sparsely populated ones. This discrepancy arises because the metrics in Table˜2 assume equal weight for all samples. WT-ASBS, designed to satisfy Requirement˜3, incorporates weighted sampling to achieve balanced efficiency across all relevant modes. Since Boltzmann populations depend exponentially on free energy, focusing on more populated regions can yield good metrics without capturing the full distribution. Therefore, we emphasize evaluations more closely aligned with the design choices discussed in the main text.
E.2 Alanine tetrapeptide
In Figure˜8, we show the distribution of pretraining data and compare the PMF obtained by reweighting (Equation˜1) for WT-ASBS, the bias (Remark˜3.1) for WTMetaD and WT-ASBS, and the reference PMF. The populations within each mode are well reproduced, while biased sampling methods (WT-ASBS and WTMetaD) capture the energetic features of the transition region between modes. This region is inaccessible even with a huge number of unbiased samples ($5\times 10^{7}$ for the reference), underscoring the necessity of biased sampling.
@@ E.2 Alanine tetrapeptide
E.3 Sampling Reactive Energy Landscapes
@@
We compare PMFs for reactions obtained with bias from WT-ASBS and WTMetaD in Figures˜9 and 10, and confirm that the final PMFs from both methods agree well. For the post-TS bifurcation, valuable chemical insights into the mechanistic understanding of reactions can be obtained from the free energy profiles. For example, while 5 is the experimentally observed product, two competing mechanisms of 3 $\to$ 4 $\to$ 5 (Paddon-Row & Warrener, 1974) vs. 3 $\to$ 5 (Houk et al., 1970) have been proposed. The 2-D PMF reveals that the initial reactive flux is biased toward product 4, but the deeper free energy basin of 5 makes it the thermodynamically favored product, consistent with previous findings that 4 forms transiently and subsequently converts to 5 (Yu et al., 2017).

### v2

@@ F.1 Alanine Dipeptide
In this section, we extend the alanine dipeptide sampling comparison from the main text to standard distributional metrics following liu2025adjoint. Specifically, we report KL divergence for 1-D marginals of torsion angles, as well as $\mathcal{W}_{2}$ and Jensen–Shannon divergence (JSD) for the joint distribution of backbone torsions $(\phi,\psi)$. Results for previous diffusion samplers using internal coordinate representations (bond distances, angles, and torsions) and for models trained in this work using Cartesian coordinates are summarized in Table˜2. We used $10^{6}$ samples to compute $D_{\text{KL}}$ and JSD, and randomly selected $10^{4}$ samples for $\mathcal{W}_{2}$ (with weighted random sampling for WT-ASBS).
@@
Table 2: Distributional metrics for alanine dipeptide. Comparison of diffusion samplers applied in internal or Cartesian coordinate representations of alanine dipeptide sampling. The KL divergence ($D_{\text{KL}}$) of five 1-D torsion angle marginals and the Wasserstein-2 distance ($\mathcal{W}_{2}$) and JSD on joint backbone torsion distribution are reported. Results from liu2025adjoint. Result from choi2025nonequilibrium. Reweighted samples are used.
@@
Figure 14: Ramachandran plots for alanine dipeptide. Comparison of backbone torsion distributions $(\phi,\psi)$ in log densities, for reference, ASBS (internal coordinates), ASBS and WT-ASBS (Cartesian coordinates), as in Table˜2. All three methods reproduce the more populated regions reasonably well, but their behavior in less populated regions differs substantially. Negligibly low-density regions are omitted for visual clarity. Taken from liu2025adjoint.
Table˜2 shows that WT-ASBS performs well on all metrics, with similar $\mathcal{W}_{2}$ performance to baseline ASBS (Cartesian) and slightly worse than reference ASBS (internal coordinates). However, the Ramachandran plots in Figure˜14 reveal that WT-ASBS better reproduces the overall density, while the other two methods either collapse into highly populated regions or leak into sparsely populated ones. This discrepancy arises because the metrics in Table˜2 assume equal weight for all samples. WT-ASBS, designed to satisfy Requirement˜3, incorporates weighted sampling to achieve balanced efficiency across all relevant modes. Since Boltzmann populations depend exponentially on free energy, focusing on more populated regions can yield good metrics without capturing the full distribution. Therefore, we emphasize evaluations more closely aligned with the design choices discussed in the main text.
F.2 Alanine Tetrapeptide
In Figure˜15, we show the distribution of pretraining data and compare the PMF obtained by reweighting (Equation˜1) for WT-ASBS, the bias (Remark˜3.1) for WTMetaD and WT-ASBS, and the reference PMF. The populations within each mode are well reproduced, while biased sampling methods (WT-ASBS and WTMetaD) capture the energetic features of the transition region between modes. This region is inaccessible even with a huge number of unbiased samples ($5\times 10^{7}$ for the reference), underscoring the necessity of biased sampling.
@@ F.2 Alanine Tetrapeptide
F.3 Sampling Reactive Energy Landscapes
@@
We compare PMFs for reactions obtained with bias from WT-ASBS and WTMetaD in Figures˜16 and 17, and confirm that the final PMFs from both methods agree well. For the post-TS bifurcation, valuable chemical insights into the mechanistic understanding of reactions can be obtained from the free energy profiles. For example, while 5 is the experimentally observed product, two competing mechanisms of 3 $\to$ 4 $\to$ 5 (paddon1974concerning) vs. 3 $\to$ 5 (houk1970novel) have been proposed. The 2-D PMF reveals that the initial reactive flux is biased toward product 4, but the deeper free energy basin of 5 makes it the thermodynamically favored product, consistent with previous findings that 4 forms transiently and subsequently converts to 5 (yu2017mechanisms).
F.4 ML Collective Variables
To illustrate the generalizability of the method to non–hand-crafted CVs, we present initial experiments coupling WT-ASBS with an ML CV learned automatically from unbiased samples of the metastable states of alanine dipeptide.
Figure 18: WT-ASBS for alanine dipeptide with ML CV. (a) CV training data from each state in the $(\phi,\psi)$ space. (b) Trained ML CV values for reference equilibrium configurations. (c) One-dimensional PMF along the Deep-TDA CV obtained with WT-ASBS and from reference unbiased configurations. (d, e) Two-dimensional PMF on $(\phi,\psi)$ from (d) WT-ASBS with reweighting and (e) reference unbiased configurations. (f) Convergence of $\Delta F$ between the two states for WT-ASBS as a function of the number of energy evaluations.
Starting from 5,000 unbiased configurations collected from each state (Figure˜18a), we trained a Deep-TDA (trizio2021enhanced) CV that maps the samples to a one-dimensional space in which each state corresponds to a single Gaussian distribution. We used the default settings of the mlcolvar package (bonati2023unified), employing a multilayer perceptron CV model that takes pairwise heavy-atom distances as inputs and includes hidden layers of sizes 24 and 12. We assign states 1 and 2 to Gaussians $\mathcal{N}(-7.0,1.0^{2})$ and $\mathcal{N}(7.0,1.0^{2})$, respectively. The resulting ML CV values for long unbiased reference configurations (including transitions) are shown in Figure˜18b: configurations from the two metastable states cluster around 7-7 and $7$, while transition configurations lie smoothly between them. We then perform WT-ASBS training using the same settings as in the main-text alanine dipeptide experiment, except for a smaller Gaussian height $h=1\times 10^{-5}$ eV to account for the reduced CV dimensionality.
The final 1-D PMF along the CV in Figure˜18c shows good agreement with the reference PMF, and $\Delta F$ approaches the reference value smoothly (Figure˜18f). The 2-D PMF on $(\phi,\psi)$ obtained by reweighting 1M samples with their bias values (Figure˜18d) also agrees reasonably well with the reference 2-D PMF (Figure˜18e), though the modes appear broadened along the $\psi$ direction, which is less well resolved by the ML CV. This behavior is consistent with the difference between baseline ASBS and WT-ASBS (Figure˜14), where WT-based weights help recover structure in sparsely sampled regions by downweighting configurations that are distinguishable in the CV space. These preliminary results suggest that WT-ASBS can be combined with ML CV pipelines to enable enhanced exploration when a good CV is not known a priori.